OpenStudy (anonymous):

(1+cos3t)/(sin3t)+(sin3t)/(1+cos3t)=2csc3t

OpenStudy (anonymous):

\[\frac{ 1+\cos3t }{ \sin3t }+\frac{ \sin3t }{ 1+\cos3t }= 2\csc3t\] what do you want? prove it?

OpenStudy (anonymous):

yes please, im having the hardest time with these questions...

OpenStudy (anonymous):

I give you idea, you do the rest. the first part \[\frac{ 1+\cos3t }{ \sin3t }= \frac{ 1 }{ \sin3t }+ \frac{ \cos3t }{ \sin3t }= \csc 3t + \cot 3t\]

OpenStudy (anonymous):

ok let me look at it

OpenStudy (anonymous):

no go, have no idea what is next

OpenStudy (anonymous):

That is pretty much it. You know what cot(x) is equal to? Also what is csc(x) equal to.

OpenStudy (anonymous):

cotx=1/tanx

OpenStudy (anonymous):

cscx=1/sinx

OpenStudy (anonymous):

what else for cotx? there should be one more

OpenStudy (anonymous):

cosx/sinx

OpenStudy (anonymous):

now look at the first step, do you see something similar to what you said

OpenStudy (anonymous):

sorry you lost me...

OpenStudy (anonymous):

you said cotx = cosx/sinx and you said cscx=1/sinx right?

OpenStudy (anonymous):

yes and that is what we converted into part of the original problem

OpenStudy (anonymous):

ok I give you mine : from yours , right side = 2/ sin3t. write it out

OpenStudy (anonymous):

now I just prove that left side * sin 2t =2, right? it is easy to do

OpenStudy (anonymous):

\[\sin 3t \left[ \frac{ 1 + cost 3t }{ \sin3t } +\frac{ \sin3t }{ 1 + \cos 3t } \right]\]

OpenStudy (anonymous):

\[= \frac{ \sin 3t + \cos 3t }{ \sin3t }+ \frac{ \sin^2 3t }{ 1 + \cos 3t }\]

OpenStudy (anonymous):

ok, how did we get from csc3t+cot3t+sin3t/1+cos3t to what you writing now?

OpenStudy (anonymous):

\[= (\frac{ sint }{ \sin3t }+ \frac{ \sin3tcos3t }{ \sin3t }+ \frac{ 1-\cos^2 3t }{1+\cos 3t })\]

OpenStudy (anonymous):

\[=1+ \cos 3t + \frac{ (1+\cos3t)(1-\cos3t) }{ 1+\cos 3t }\]

OpenStudy (anonymous):

and then you have 1+1 =2. the proof done. is it clear?

OpenStudy (anonymous):

I make the right side2 csc 3t = 2/ sin3t. and then time both sides to sin 3t

OpenStudy (anonymous):

do you get mine? think of it. write out and it will make sense to you.

OpenStudy (anonymous):

ok let me look at it and see, im still not sure why we did the very first step as we ended up working out the original set up

OpenStudy (anonymous):

our goal is prove the original equation is true, that means you must prove the left side = right side. some equation is not true at all, for example : how can you prove 1+1 =3? fortunately, your problem is prove it true, that means it is true already, just explain in which logic, it is true. done

OpenStudy (anonymous):

ok, let me work it out on paper and see what I come up with. Thanks for your help.

OpenStudy (anonymous):

you are welcome